OpenStudy (anonymous):

pleas help someone, step by step= 5x+y+3=-14=8y-x x+2x=y+4 substitution method

OpenStudy (anonymous):

In the substitution we want to either express x in terms of y, or y in terms of x so that we can get numerical answers. So... lets begin! \[5x+y+3=-14-8y-x\] \[x+2x=y+4\] I'm going to first start by simplifying the second equation: \[x+2x=y+4\] \[3x=y+4\] (x+2x is equal to 3x) \[3x-4=y\] (moving 4 to the right side by subtracting 4 from both sides) So, now we know that y is equal to the value of 3x's minus four, so lets plug in 3x-4 in for every "y" in the first equation

OpenStudy (anonymous):

Knowing \[y=3x-4\] I'm going to plug that in for every "y" in the first equation. So... let's do that. ----- \[5x+y+3=-14-8y-x\]| Original equation \[5x+(3x-4)+3=-14-(8(3x-4))-x\] Substituted 3x-4 for every "y" in the equation \[5x+(3x-4)+3=-14-(24x-32)-x\] Simplified 8 times 3x-4 \[5x+3x-4+3=-14-24x-32-x\] Removed parenthesis \[8x-1=-14-24x-32-x\] 5x+3x is 8x, 3-4 is -1 (so more simplification) \[8x-1=-46-23x\] -14-32 is -46, 24x-x is 23x (so I simplified the right hand side)

OpenStudy (anonymous):

\[8x-1=-46-23x\] (Recopied equation from last post) \[8x=-45-23x\] Added 1 to both sides \[31x=-45\] Added 23x to both sides \[x=\frac{-45}{31}\] Divide both sides by 31 to get x by its self.

OpenStudy (anonymous):

So now we know that \[x=\frac{-45}{31}\] Let's revisit the equation we simplified in the very beginning: \[y=3x-4\] Now that we know what the value of x is, we just simply plug in that value into the equation and solve for y!

OpenStudy (anonymous):

ok, p

OpenStudy (anonymous):

im working on it

OpenStudy (anonymous):

y= y=134/31-4

OpenStudy (anonymous):

Yep. so: \[y=\frac{131}{31}-4\] \[x=-\frac{45}{31}\]

OpenStudy (anonymous):

so is that the final answer